OpenStudy (anonymous):

sum from n=0 ->infinity (1+ sin(n))/ (10^n)

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (1+ \sin(n))/(10^n)\]

OpenStudy (anonymous):

convergent...divergent?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \sin(n)/10^n\] to use as the comparison test

OpenStudy (anonymous):

what happened to the sine?

OpenStudy (anonymous):

\[\sin (n) \le 1\]

OpenStudy (anonymous):

so we can ignore it when checking for convergence?

OpenStudy (anonymous):

you have to solve it exactly by Firier series,Is the problem in that area?

OpenStudy (anonymous):

no i just have to determine if it is convergent or divergent by using the comparison test

OpenStudy (anonymous):

so mukushla solved it and showed it is convergence.

OpenStudy (anonymous):

i didn't quit understand how we went (1+sin(n))/(10n) to 2/(10n) to (2n)/(10n)

OpenStudy (anonymous):

bot he had a mistake and should put 2 instead 2^n

OpenStudy (anonymous):

but whyyyyy?

OpenStudy (anonymous):